# A2Seek: Towards Reasoning-Centric Benchmark for Aerial Anomaly Understanding

### v1

@@ 4 Methodology
Anomaly detection from aerial perspectives poses unique challenges, requiring both precise spatial localization (“Where is the anomaly?”) and comprehensive semantic reasoning (“Why is it an anomaly?”). To address these challenges, we propose A2Seek-R1, a reasoning-centric framework specifically designed for aerial anomaly understanding. The framework integrates two key stages: (1) Supervised Fine-Tuning (SFT), which leverages Graph-of-Thought (GoT) annotations to activate the model’s initial reasoning capabilities by structuring reasoning into explicit semantic stages, and (2) Reinforcement Fine-Tuning (RFT), which employs Aerial Group Relative Policy Optimization (A-GRPO) to refine the model’s reasoning and localization strategies through task-specific reward functions. Among them, the seeking reward enables the model to dynamically identify regions requiring additional high-resolution context for further analysis, while the length reward encourages concise answers in simple scenarios and allocates more reasoning effort to complex and challenging scenes. Thought A-GRPO-driven reinforcement fine-tuning, the model not only achieves accurate anomaly detection but also generates interpretable and verifiable reasoning traces, making it effective for real-world aerial applications.
@@ 5.1 Implementation Details
We employ Qwen2.5VL-3B-Instruct bai2025qwen2 as the base model due to its strong performance in vision-language understanding. We employed LoRA hu2022lora to adjust all fully connected layers within the model, and set a learning rate of $1\mathrm{e}{-5}$, using 4 epochs for SFT and 1 epoch for RL. Input resolution is set to $896\times 448$, with $4$ frames per sequence. Batch size is 1, and gradient accumulation is performed over 16 steps. A cosine learning rate scheduler with a warm-up phase (5%) is used. All experiments are conducted in PyTorch on a platform with an Intel Xeon Platinum 8350 CPU, four NVIDIA A100 GPUs, and 1,024 GB of memory.
@@
Method Scene Indices Average S00 S01 S02 S03 S04 S05 S06 S07 S08 S09 SSRL li2022scale 11.50 10.20 08.30 09.40 20.10 12.70 05.80 14.20 16.80 11.00 12.00 AnoamlyRuler yang2024anomalyruler 17.40 15.00 09.80 22.50 25.60 14.90 06.00 19.30 21.70 18.80 17.10 LAVAD zanella2024harnessing 12.80 15.20 11.00 19.50 24.70 20.50 04.30 19.10 22.10 14.90 16.41 Holmes-VAU zhang2024holmes 17.00 18.50 16.00 23.40 25.00 20.50 13.34 21.12 27.20 17.60 19.97 LLavaVideo liu2024visual 12.00 10.00 16.00 15.00 24.00 13.00 04.00 23.00 12.00 22.00 15.10 InternVL-3.0 chen2024internvl 35.62 49.75 48.21 47.08 27.43 23.94 72.40 29.64 36.99 42.33 41.34 Zero-shot bai2025qwen2 44.83 44.92 35.15 23.88 22.42 19.16 62.80 30.05 29.44 27.62 34.03 RL-Zero 00.56 14.32 07.59 02.06 29.10 04.94 01.60 16.07 18.84 29.10 12.42 ANS-SFT 07.87 32.10 45.89 25.12 32.63 16.79 44.00 34.29 47.95 14.29 30.09 CoT-SFT 42.81 51.36 42.41 62.20 23.47 25.02 73.60 25.36 24.66 39.68 41.06 GoT-SFT 49.44 54.32 54.02 64.96 32.34 23.20 63.20 26.43 54.45 41.80 46.42 A2Seek-R1 51.12 56.54 55.36 75.60 40.84 26.86 81.60 38.21 61.30 43.92 53.14
@@ 5.2 Overall Performance
To highlight the contributions of our dataset and method, we design the following experimental settings: (1) Zero-Shot, serving as a baseline without any fine-tuning; (2) RL-Zero, which applies GRPO with basic format and accuracy rewards; (3) ANS-SFT, utilizing anomaly labels for supervised fine-tuning; (4) CoT-SFT, incorporating chain-of-thought to guide inference; and (5) GoT-SFT, leveraging our graph-of-thought reasoning data to enable structured anomaly understanding.
@@
(a) Average precision and language metrics.
Method $AP_{c}$ $BLEU$ $METEOR$ $ROUGE$ $CIDEr$ Zero-Shot 34.03 0.2835 0.2145 0.3263 0.8147 ANS-SFT 31.10 0.3821 0.3452 0.4379 1.0462 GOT-SFT 46.42 0.4478 0.3325 0.4439 1.1528 A2Seek-R1 53.14 0.4564 0.3543 0.4882 1.2989
(b) Localization performance.
@@
Method $mIoU$ $AP_{0.00}$ $AP_{0.25}$ $AP_{0.50}$ ${AP_{0.75}}$ ${AP_{0.90}}$ Zero-Shot 03.50 72.40 03.94 01.81 00.63 0.00 ANS-SFT 17.05 50.70 40.01 31.94 15.74 0.70 GOT-SFT 20.81 43.85 35.88 27.60 16.01 1.03 A2Seek-R1 26.03 53.31 45.43 35.33 20.11 4.34
@@
Language Semantic Evaluation. To evaluate the ability of models to handle semantic ambiguity and similarity in aerial anomaly understanding, we adopt language-based metrics such as BLEU papineni2002bleu , METEOR banerjee2005meteor , ROUGE lin2004rouge , and CIDEr vedantam2015cider . These metrics capture both lexical and structural alignment between predicted and ground-truth descriptions. As shown in Table 3(a), A2Seek-R1 achieves the highest scores across all metrics, with BLEU of 0.4564, METEOR of 0.3543, ROUGE of 0.4882, and CIDEr of 1.2989. These results demonstrate its ability to generate accurate and semantically meaningful descriptions even under ambiguous scenarios. By leveraging visual evidence during reasoning, A2Seek ensures that its language-based predictions are contextually grounded, effectively bridging the gap between semantic understanding and visual perception.
@@
Supporting Region Grounding Accuracy. To further evaluate whether the model relies on visual information to understand anomalies, we assess its localization performance using both mIoU and AP under varying IoU thresholds. As shown in Table 3(b), A2Seek-R1 achieves the highest mIoU score of 26.03%, outperforming Zero-Shot and GoT-SFT by 22.53% and 5.22%, respectively. Notably, although Zero-Shot and ANS-SFT achieve relatively high AP values in $AP_{0.0}$, this is mainly attributed to their tendency to predict overly broad and vague regions, rather than accurately indicating the anomalous areas. This limitation is further reflected in their significantly degraded performance (nearly zero in $AP_{0.90}$) at higher IoU thresholds, indicating a tendency toward random guessing. These results confirm that the spatial reasoning mechanism introduced in our approach enhances the model’s ability to precisely localize subtle and ambiguous anomalies, thereby improving both detection accuracy and interpretability for real-world aerial anomaly understanding.
(a) Overall accuracy.
(b) Data efficiency.
(c) OOD generalization.
Figure 3: Performance comparison of different setting on A2Seek benchmark.
@@
Data Efficiency and Generalization. The in-domain (ID) data in A2Seek includes scenarios present in the training set, while the out-of-domain (OOD) data comprises unseen scenarios, specifically designed to assess the model’s generalization capability in novel and challenging settings. As shown in Figure 3, A2Seek-R1 demonstrates both high data efficiency on in-domain data and strong generalization to out-of-domain scenarios. On ID data (Figure 3(b)), A2Seek-R1 achieves consistent performance improvements throughout training, whereas GoT-SFT shows stagnation or minor fluctuations. Notably, on OOD data (Figure 3(c)), A2Seek-R1 continues to gain accuracy as training progresses, reaching 48.98% at 400 steps, a relative improvement of 4.57% percentage points. In contrast, GoT-SFT degrades from 44.41% to 41.84% over the same period. This observation suggests that A2Seek-R1 not only leverages training data more effectively but also exhibits better robustness under domain shifts.
@@ 5.3 Ablation Studies
Table 4 presents the results of ablation studies on the A2Seek dataset, analyzing the contributions of supervised fine-tuning (SFT) strategies and reinforcement learning (RL) reward components to anomaly detection ($AP_{c}$) and localization ($mIoU$) performance.
Table 4: Ablation studies of the A2Seek-R1 framework.
Idx Supervised Fine-Tuning Reinforcement Learning Metric ANS CoT GoT Accuracy Reward Location Reward Seeking Reward $AP_{c}$ $mIoU$ 0 31.10 12.13 1 38.24 17.05 2 46.42 20.81 3 12.42 09.66 4 52.82 18.77 5 51.78 24.03 6 53.14 26.03
@@ 6 Conclusions
Summary. This work introduces A2Seek, a large-scale, reasoning-centric dataset for aerial anomaly understanding. Built upon diverse aerial scenarios, A2Seek is meticulously annotated with fine-grained anomaly labels, spatiotemporal bounding boxes, and structured reasoning graphs. These annotations enable models to not only detect anomalies but also provide interpretable explanations for “why” and “where” anomalies occur. Building on this benchmark, we introduce A2Seek-R1, a novel multi-stage reinforcement fine-tuning framework that integrates graph-of-thought (GoT)-guided supervised fine-tuning, and an improved A-GRPO algorithm with tailored reward functions, including a seeking mechanism that simulates UAV flight behavior. This framework achieves state-of-the-art performance in both accuracy and interpretability, excelling in complex anomaly scenarios and demonstrating strong cross-domain generalization capabilities.
Limitations and Societal Impact. Despite strong performance, the complexity of reasoning chains and strategy learning limits inference speed and deployment efficiency. Additionally, given its potential applications in sensitive areas like public safety, privacy protection and algorithmic fairness must be prioritized to avoid societal risks from misjudgments.
@@ A.1 Dataset Acquisition Protocol
To address the challenges of spatial localization and scene generalization, the data collection process was carefully designed to maximize diversity and realism. The A2Seek dataset was collected using a DJI M30T drone equipped with wide-angle, telephoto, and infrared cameras. The drone operated at varying altitudes (10 to 60 meters) and speeds (0 to 20 m/s) to capture a wide range of scale variations for anomalous objects. Flight trajectories included hovering, linear cruising, curved circling, and area scanning, reflecting dynamic viewpoint changes. The dataset comprises 542 untrimmed 4K videos recorded at 30 FPS across 10 campus environments, subdivided into 27 sub-scenes. Each video has an average duration of 153.74 seconds, with a total duration exceeding 23 hours. Among these, 19.3 hours feature normal behaviors, while 3.79 hours capture diverse and complex anomalies, including parallel, sequential, occluded, and scene-dependent events. As shown in Figure 5, our dataset includes over 20 anomaly categories, such as falling, fighting, playing with water, and jaywalking. Spanning nearly a year of recording, it covers various seasons, times of day (daytime, nighttime, twilight), and environmental conditions, including clear, cloudy, foggy, and rainy weather. This diversity provides a robust foundation for developing and evaluating aerial anomaly detection models.
Table 5 presents a unified overview of all categorical codes used in this paper. The Type column groups related codes, the Code column lists shorthand labels, and the Definition column provides detailed descriptions.
@@
Type Code Definition Scene S00 Miscellaneous: other or unspecified scenes S01 Roadway: street and vehicular paths S02 Sidewalk: pedestrian walkways along roads S03 Playground: open recreational areas S04 Open Area: unstructured outdoor space S05 Park: landscaped public green spaces S06 Rooftop: building rooftop areas S07 Entrance: school gate or entryway region S08 Wall Zone: perimeter walls and boundaries S09 Academic Building: building facade and courtyard Action E00 Normal: no anomalous activity E01 Loitering: prolonged stationary presence E02 Trespassing on Lawn: entering grass areas illegally E03 Running: fast movement on foot E04 Animal: wildlife appearing in view E05 Vandalism: intentional damage to property E06 Falling: loss of balance and fall incident E07 Unable to Stand: person remains on ground after fall E08 Playing with Water: interacting with water bodies E09 Unconventional Vehicle: use of nonstandard transport E10 Wall Climbing: scaling vertical surfaces E11 Carrying Weapon: person holds or draws weapon E12 Jaywalking: crossing road outside crosswalk E13 Bicycling: riding a bicycle on pathways E14 Bullying: aggressive behavior towards others E15 Lost Item: dropping or forgetting personal items E16 Sneaking: moving stealthily or furtively E17 Theft: unauthorized taking of property E18 Fighting: physical aggression between persons E19 Robbery: theft involving force or threat E20 Littering: discarding waste improperly Height H0 10–20 m (Low altitude flight) H1 20–50 m (High altitude flight) Velocity M0 0–3 m/s (Low speed) M1 3–10 m/s (Moderate speed) M2 10–20 m/s (High speed) Time of Day L0 10:00–16:00 (Daytime) L1 08:00–10:00, 16:00–18:00 (Morning/Evening) L2 18:00–22:00 (Evening/Night) Weather W0 Clear: no precipitation or clouds W1 Cloudy: partial cloud cover W2 Overcast: full cloud cover W3 Rain: rainfall conditions W5 Fog: reduced visibility due to fog W8 Night: after sunset until pre-dawn Risk Level High-risk Actions likely to cause serious harm or property damage (E05, E11, E14, E17, E18, E19) Moderate-risk Actions that may result in moderate safety concerns or disruption (E06, E07, E09, E10, E12, E16) Low-risk Actions considered minor with minimal safety risk (E00, E01, E02, E03, E04, E08, E13, E15, E20)
Our collection workflow for A2Seek condenses a campus-scene drone shoot into a single, reproducible loop that still preserves fine-grained control:
  • Scene–Time Grid: ten functional zones, further split into 27 sub-scenes, are revisited across four seasons, three day parts (dawn 06:00 to 08:00, daylight 10:00 to 16:00, night 19:00 to 22:00) and five weather classes, producing 455 UHD clips.
  • Airframe & Optics: a DJI M30T (1-inch CMOS, 20 MP, 3840$\times$2160@30fps) plus a co-aligned 640$\times$512 thermal core for nocturnal sorties; pose is logged at 50 Hz by the on-board GNSS–IMU.
  • Flight Envelope: grid missions with 80 % side / 70 % forward overlap (routine coverage) that can be hot-switched to follow/orbit mode for rare anomalies; altitude 10-60 m above ground level ($\leq$ 30 m at night) and ground speed 0-15 m/s.
  • Data Handling: RGB, thermal and telemetry are streamed to a 512 GB UHS-II SD card, then mirrored to an offline workstation on landing; filenames encode UTC timestamps and scene IDs for instant indexing.
  • Quality Gate: each day, 10% of sorties are spot-checked by two senior annotators for focus, exposure and occlusion; flights with privacy inspection or other unqualified situation are queued for reshoot.
This lean yet rigorous protocol furnishes high-resolution, context-rich footage tailored to drone-view anomaly detection while keeping field operations swift and fully repeatable.

### v2

@@ 4 Methodology
Anomaly detection from aerial perspectives poses unique challenges, requiring both precise spatial localization (“Where is the anomaly?”) and comprehensive semantic reasoning (“Why is it an anomaly?”). To address these challenges, we propose A2Seek-R1, a reasoning-centric framework specifically designed for aerial anomaly understanding. The framework integrates two key stages: (1) Supervised Fine-Tuning (SFT), which leverages Graph-of-Thought (GoT) annotations to activate the model’s initial reasoning capabilities by structuring reasoning into explicit semantic stages, and (2) Reinforcement Fine-Tuning (RFT), which employs Aerial Group Relative Policy Optimization (A-GRPO) to refine the model’s reasoning and localization strategies through task-specific reward functions. Among them, the seeking reward enables the model to dynamically identify regions requiring additional high-resolution context for further analysis, while the length reward encourages concise answers in simple scenarios and allocates more reasoning effort to complex and challenging scenes. Through A-GRPO-driven reinforcement fine-tuning, the model not only achieves accurate anomaly detection but also generates interpretable and verifiable reasoning traces, making it effective for real-world aerial applications.
@@ 5.1 Implementation Details
We employ Qwen2.5-VL-3B-Instruct bai2025qwen2 as the base model due to its strong performance in vision-language understanding. We employed LoRA hu2022lora to adjust all fully connected layers within the model, and set a learning rate of $1\mathrm{e}{-5}$, using 4 epochs for SFT and 1 epoch for RL. Input resolution is set to $896\times 448$, with $4$ frames per sequence. Batch size is 1, and gradient accumulation is performed over 16 steps. A cosine learning rate scheduler with a warm-up phase (5%) is used. All experiments are conducted in PyTorch on a platform with an Intel Xeon Platinum 8350 CPU, four NVIDIA A100 GPUs, and 1,024 GB of memory.
@@
Method Scene Indices Average S00 S01 S02 S03 S04 S05 S06 S07 S08 S09 SSRL li2022scale 11.50 10.20 08.30 09.40 20.10 12.70 05.80 14.20 16.80 11.00 12.00 HSTforU le2025hstforu 47.41 50.18 43.52 30.03 29.68 25.01 09.70 37.28 48.72 33.69 47.66 ANDT jin2022anomaly 10.89 39.00 48.82 25.21 20.20 19.94 30.18 37.68 49.99 28.63 40.42 AnomalyRuler yang2024anomalyruler 17.40 15.00 09.80 22.50 25.60 14.90 06.00 19.30 21.70 18.80 17.10 LAVAD zanella2024harnessing 12.80 15.20 11.00 19.50 24.70 20.50 04.30 19.10 22.10 14.90 16.41 Holmes-VAU zhang2024holmes 17.00 18.50 16.00 23.40 25.00 20.50 13.34 21.12 27.20 17.60 19.97 LLavaVideo liu2024visual 12.00 10.00 16.00 15.00 24.00 13.00 04.00 23.00 12.00 22.00 15.10 InternVL-3.0 chen2024internvl 35.62 49.75 48.21 47.08 27.43 23.94 72.40 29.64 36.99 42.33 41.34 Zero-shot bai2025qwen2 44.83 44.92 35.15 23.88 22.42 19.16 62.80 30.05 29.44 27.62 34.03 RL-Zero 00.56 14.32 07.59 02.06 29.10 04.94 01.60 16.07 18.84 29.10 12.42 ANS-SFT 07.87 32.10 45.89 25.12 32.63 16.79 44.00 34.29 47.95 14.29 30.09 CoT-SFT 42.81 51.36 42.41 62.20 23.47 25.02 73.60 25.36 24.66 39.68 41.06 GoT-SFT 49.44 54.32 54.02 64.96 32.34 23.20 63.20 26.43 54.45 41.80 46.42 A2Seek-R1 51.12 56.54 55.36 75.60 40.84 26.86 81.60 38.21 61.30 43.92 53.14
@@ 5.2 Overall Performance
To highlight the contributions of our dataset and method, we design the following experimental settings: (1) Zero-Shot, serving as a baseline without any fine-tuning; (2) RL-Zero, which applies GRPO with basic format and accuracy rewards; (3) ANS-SFT, utilizing anomaly labels for supervised fine-tuning; (4) CoT-SFT, incorporating chain-of-thought to guide inference; and (5) GoT-SFT, leveraging our graph-of-thought reasoning data to enable structured anomaly understanding. We evaluate all methods using category-level average precision ($AP_{c}$) for anomaly detection and mean Intersection over Union ($mIoU$) for localization, providing a balanced assessment of classification and grounding performance.
@@
(a) Average precision and language metrics
@@
Method $AP_{c}$ BLEUBLEU METEORMETEOR ROUGEROUGE CIDErCIDEr $Samples/s$ Zero-Shot 34.03 0.2835 0.2145 0.3263 0.8147 0.6097 ANS-SFT 31.10 0.3821 0.3452 0.4379 1.0462 0.9174 GOT-SFT 46.42 0.4478 0.3325 0.4439 1.1528 0.2890 A2Seek-R1 53.14 0.4564 0.3543 0.4882 1.2989 0.3267
(b) Localization performance
Method mIoUmIoU $AP_{0.00}$ $AP_{0.25}$ $AP_{0.50}$ $AP_{0.75}$} $AP_{0.90}$} Zero-Shot 03.50 72.40 03.94 01.81 00.63 0.00 ANS-SFT 17.05 50.70 40.01 31.94 15.74 0.70 GOT-SFT 20.81 43.85 35.88 27.60 16.01 1.03 A2Seek-R1 26.03 53.31 45.43 35.33 20.11 4.34
@@
Language Semantic Evaluation. To evaluate the ability of models to handle semantic ambiguity and similarity in aerial anomaly understanding, we adopt language-based metrics such as BLEU papineni2002bleu, METEOR banerjee2005meteor, ROUGE lin2004rouge, and CIDEr vedantam2015cider. These metrics capture both lexical and structural alignment between predicted and ground-truth descriptions. As shown in Table 3(a), A2Seek-R1 achieves the highest scores across all metrics, with BLEU of 0.4564, METEOR of 0.3543, ROUGE of 0.4882, and CIDEr of 1.2989. The throughput is 0.3267 samples per second, which is lower than ANS-SFT (0.9174) and Zero-Shot (0.6097) because our method allocates a larger token budget at test time for reasoning, leading to better anomaly understanding. These results demonstrate its ability to generate accurate and semantically meaningful descriptions even under ambiguous scenarios. By leveraging visual evidence during reasoning, A2Seek-R1 ensures that its language-based predictions are contextually grounded, effectively bridging the gap between semantic understanding and visual perception.
@@
Supporting Region Grounding Accuracy. To further evaluate whether the model relies on visual information to understand anomalies, we assess its localization performance using both mIoU and AP under varying IoU thresholds. As shown in Table 3(b), A2Seek-R1 achieves the highest mIoU score of 26.03%, outperforming Zero-Shot and GoT-SFT by 22.53% and 5.22%, respectively. Notably, although Zero-Shot and ANS-SFT achieve relatively high AP values in $AP_{0.00}$, this is mainly attributed to their tendency to predict overly broad and vague regions, rather than accurately indicating the anomalous areas. This limitation is further reflected in their significantly degraded performance (nearly zero in $AP_{0.90}$) at higher IoU thresholds, indicating a tendency toward random guessing. These results confirm that the spatial reasoning mechanism introduced in our approach enhances the model’s ability to precisely localize subtle and ambiguous anomalies, thereby improving both detection accuracy and interpretability for real-world aerial anomaly understanding.
(a) Overall accuracy
(b) Data efficiency
(c) OOD generalization
Figure 3: Performance comparison of different settings on A2Seek benchmark.
@@
Data Efficiency and Generalization. The in-domain (ID) data in A2Seek includes scenarios present in the training set, while the out-of-domain (OOD) data comprises unseen scenarios, specifically designed to assess the model’s generalization capability in novel and challenging settings. As shown in Figure 3, A2Seek-R1 demonstrates both high data efficiency on in-domain data and strong generalization to out-of-domain scenarios. On ID data (Figure 3(b)), A2Seek-R1 achieves consistent performance improvements throughout training, whereas GoT-SFT shows stagnation or minor fluctuations. Notably, on OOD data (Figure 3(c)), A2Seek-R1 continues to gain accuracy as training progresses, reaching 48.98% at 400 steps, a relative improvement of 4.57%. In contrast, GoT-SFT degrades from 44.41% to 41.84% over the same period. This observation suggests that A2Seek-R1 not only leverages training data more effectively but also exhibits better robustness under domain shifts.
Table 4: Ablation studies of the A2Seek-R1 framework.
Idx Supervised Fine-Tuning Reinforcement Learning Metric ANS CoT GoT Accuracy Reward Location Reward Seeking Reward $AP_{c}$ mIoUmIoU 0 31.10 12.13 1 38.24 17.05 2 46.42 20.81 3 12.42 09.66 4 52.82 18.77 5 51.78 24.03 6 53.14 26.03
@@ 5.3 Ablation Studies
Table 4 presents the results of ablation studies on the A2Seek dataset, analyzing the contributions of supervised fine-tuning (SFT) strategies and reinforcement learning (RL) reward components.
@@
Table 5: Effectiveness of A2Seek-R1 on different base models.
[…]
Impact of Foundation Models. As shown in Table 5, we evaluate the generalization of A2Seek-R1 on two additional vision-language models beyond the base model, i.e., QwenVL-2.5-3B-Instruct bai2025qwen2. On QwenVL-2.5-7B-Instruct bai2025qwen2, it improves AP by 12.42% and mIoU by 21.47%, indicating effective scaling with capacity. On InternVL-3.0-2B-Instruct chen2024internvl, it increases AP from 20.03% to 48.76% and mIoU from 0.10% to 18.89%. These results demonstrate consistent gains across model sizes and architectures, supporting strong generalization and transferability.
@@ 6 Conclusions
This work introduces A2Seek, a large-scale, reasoning-centric dataset for aerial anomaly understanding. Built upon diverse aerial scenarios, A2Seek is meticulously annotated with fine-grained anomaly labels, spatiotemporal bounding boxes, and structured reasoning graphs. These annotations enable models to not only detect anomalies but also provide interpretable explanations for “why” and “where” anomalies occur. Building on this benchmark, we introduce A2Seek-R1, a novel multi-stage reinforcement fine-tuning framework that integrates graph-of-thought (GoT)-guided supervised fine-tuning, and an improved A-GRPO algorithm with tailored reward functions, including a seeking mechanism that simulates UAV flight behavior. This framework achieves state-of-the-art performance in both accuracy and interpretability, excelling in complex anomaly scenarios and demonstrating strong cross-domain generalization capabilities.
Limitations and Societal Impact. Despite strong performance, the complexity of reasoning chains and strategy learning limits inference speed and deployment efficiency. The framework mainly focuses on spatial-level anomaly perception, whereas the understanding and modeling of long-duration temporal anomalies are still limited, which we regard as an important avenue for future exploration. Additionally, given its potential applications in sensitive areas like public safety, privacy protection and algorithmic fairness must be prioritized to avoid societal risks from misjudgments.
Acknowledgments and Disclosure of Funding.
This work was supported in part by the Science and Technology Innovation Key R&D Program of Chongqing under Grant No. CSTB2023TIAD-STX0016, in part by the National Natural Science Foundation of China under Grants No. 62472060, U23A20318, and 62221005, in part by the Natural Science Foundation of Chongqing under Grants No. CSTB2024NSCQ-QCXMX0060 and CSTB2023NSCQ-LZX0061, in part by the Science and Technology Research Program of Chongqing Municipal Education Commission under Grant No. KJZD-K202300604, in part by the Chongqing Doctoral Student Innovation Talent Project under Grant No. CYB240241, in part by the Chongqing University of Posts and Telecommunications Ph.D. Innovative Talents Project under Grant No. BYJS202404, and in part by the Chongqing Institute for Brain and Intelligence.
@@ A.1 Dataset Acquisition Protocol
To address the challenges of spatial localization and scene generalization, the data collection process was carefully designed to maximize diversity and realism. The A2Seek dataset was collected using a DJI M30T drone equipped with wide-angle, telephoto, and infrared cameras. The drone operated at varying altitudes (10 to 60 meters) and speeds (0 to 20 m/s) to capture a wide range of scale variations for anomalous objects. Flight trajectories included hovering, linear cruising, curved circling, and area scanning, reflecting dynamic viewpoint changes. The dataset comprises 542 untrimmed 4K videos recorded at 30 FPS across 10 campus environments, subdivided into 27 sub-scenes. Each video has an average duration of 153.74 seconds, with a total duration exceeding 23 hours. Among these, 19.3 hours feature normal behaviors, while 3.79 hours capture diverse and complex anomalies, including parallel, sequential, occluded, and scene-dependent events. As shown in Figure 5, our dataset includes over 20 anomaly categories, such as falling, fighting, playing with water, and jaywalking. Spanning nearly a year of recording, it covers various seasons, times of day (daytime, nighttime, twilight), and environmental conditions, including clear, cloudy, foggy, and rainy weather (as shown in Figure 6). This diversity provides a robust foundation for developing and evaluating aerial anomaly detection models. Table 6 presents a unified overview of all categorical codes used in this paper. The Type column groups related codes, the Code column lists shorthand labels, and the Definition column provides detailed descriptions.
Figure 6: Examples of scenes from the A2Seek dataset.
Our collection workflow for A2Seek streamlines a campus drone survey into a single, reproducible loop while preserving fine-grained control:
  • Scene & Time: ten functional zones, further divided into 27 sub-scenes, are revisited across four seasons, three day parts (twilight: 06:00 to 08:00 or 16:00 to 19:00, daylight: 10:00 to 16:00, night: 19:00 to 22:00) and five weather classes, producing initial multimodal videos.
[…]
  • Flight Envelope: grid missions with 80% side and 70% forward overlap (for routine coverage), dynamically switchable to follow or orbit mode for rare anomalies, at altitudes of 10–60 m above ground level ($\leq$ 30 m at night) and ground speeds of 0–15 m/s.
  • Data Handling: RGB, thermal and telemetry streams are recorded to a 512 GB UHS-II SD card, then mirrored to an offline workstation upon landing; filenames are encoded with UTC timestamps and scene IDs for instant indexing.
  • Quality Gate: all sorties are checked daily by two senior annotators for focus, exposure and occlusion. Flights flagged for privacy concerns or other disqualifying conditions are scheduled to be re-flown.
This lean yet rigorous protocol provides high-resolution, context-rich footage tailored to drone-view anomaly detection while keeping field operations swift and fully repeatable.
Figure 7: Four–stage annotation workflow. Step 1 (blue) Temporal annotators use our in-house GUI to mark the start/end frames and class of every anomalous episode, exporting a JSON timeline. Step 2 (salmon) For the first frame of each event, experts draw a bounding box around the anomalous region and supply a natural-language description, yielding single-frame spatial seeds. Step 3 (green) A pretrained tracker propagates each seed through the clip to form full-length trajectories; an automated checker screens the results, flags uncertain cases for human correction, and funnels all approved tracks into the spatial-label repository. Step 4 (violet) Vision-language models (VLMs) ingest the temporal tags, spatial tracks, and human captions; via chain-of-thought reasoning, they merge these cues into consolidated frame-level annotations, producing the final label set.
@@
Type Code Definition Scene S00 Miscellaneous: other or unspecified scenes S01 Roadway: street and vehicular paths S02 Sidewalk: pedestrian walkways along roads S03 Playground: open recreational areas S04 Open Area: unstructured outdoor space S05 Park: landscaped public green spaces S06 Rooftop: building rooftop areas S07 Entrance: school gate or entryway region S08 Wall Zone: perimeter walls and boundaries S09 Academic Building: building facade and courtyard Action E00 Normal: no anomalous activity E01 Loitering: prolonged stationary presence near entrances or perimeter areas E02 Trespassing on Lawn: entering grass areas illegally E03 Running: fast movement on foot in restricted or non-sport areas E04 Animal: appearance of animals within the scene E05 Vandalism: intentional damage to property E06 Falling: loss of balance leading to a fall E07 Unable to Stand: person unable to recover after a fall E08 Playing with Water: interacting with fountains or water bodies E09 Unconventional Vehicle: use of nonstandard or unauthorized vehicles E10 Wall Climbing: scaling vertical surfaces E11 Carrying Weapon: person holding or drawing a weapon E12 Jaywalking: crossing road outside crosswalk E13 Bicycling: riding a bicycle in pedestrian or restricted areas E14 Bullying: aggressive behavior towards others E15 Lost Item: dropping or leaving behind personal belongings E16 Sneaking: moving stealthily or furtively E17 Theft: unauthorized taking of property E18 Fighting: physical aggression between persons E19 Robbery: theft involving force or threat E20 Littering: discarding waste improperly in public areas Height H0 10–20 m (Low altitude flight) H1 20–50 m (High altitude flight) Velocity M0 0–3 m/s (Low speed) M1 3–10 m/s (Moderate speed) M2 10–20 m/s (High speed) Time of Day L0 10:00–16:00 (Daytime) L1 08:00–10:00, 16:00–18:00 (Morning/Evening) L2 18:00–22:00 (Evening/Night) Weather W0 Clear: no precipitation or clouds W1 Cloudy: partial cloud cover W2 Overcast: full cloud cover W3 Rain: rainfall conditions W5 Fog: reduced visibility due to fog W8 Night: after sunset until pre-dawn Risk Level High-risk Actions likely to cause serious harm or property damage (E05, E11, E14, E17, E18, E19) Moderate-risk Actions that may result in moderate safety concerns or disruption (E06, E07, E09, E10, E12, E16) Low-risk Actions considered minor with minimal safety risk (E00, E01, E02, E03, E04, E08, E13, E15, E20)